An aircraft is moving horizontally with a speed of 60 m/s. At the height of 2 km, an object is dropped from the aircraft. Acceleration due to gravity = 9.8 m/s2, what is the time interval before the object hits the ground.

1
Expert's answer
2022-05-05T13:56:17-0400
"t=\sqrt{\frac{2H}{g}}"

"t=\sqrt{\frac{2*2000\:\rm m}{9.8\;\rm m/s^2}}=20\:\rm s"